OpenStudy (anonymous):

Prove the following vector identities letting: V = u i + v j + w k and ∇= ∂╱∂x i + ∂╱∂y j + ∂╱∂z k that: ∇∙(fV̅) = f (∇∙V̅) + V̅ ∙∇f

OpenStudy (anonymous):

What is fV ?

OpenStudy (anonymous):

f is a scalar

OpenStudy (anonymous):

v is the vector

OpenStudy (anonymous):

I am aware. I meant write down what it actually is.

OpenStudy (anonymous):

Explicitly

OpenStudy (anonymous):

f (u)i + f(v) j + f(w)k

OpenStudy (anonymous):

f is multiply u, v, and w, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So what's the divergence of that?

OpenStudy (anonymous):

∂╱∂x i + ∂╱∂y j + ∂╱∂z k

OpenStudy (anonymous):

No, what is the divergence of the expression you wrote for fV ?

OpenStudy (anonymous):

i'm not quite following what your asking me

OpenStudy (anonymous):

\[f\vec{V} = fu\vec{i} + fv\vec{j} + fw\vec{k} \] what is \[\vec{\nabla} \cdot (f\vec{V}) \] ?

OpenStudy (anonymous):

d f(u)/dx i + d f(v)/dy j + d f(w)/dz k

OpenStudy (anonymous):

the i, j, and k are gone.

OpenStudy (anonymous):

Okay, so expand that out.

OpenStudy (anonymous):

so just d f(u)/dx + d f(v)/dy + d f(w)/dz

OpenStudy (anonymous):

That's right. Now show that that equals the identity that you were given above.

OpenStudy (anonymous):

well the right hand side of the equation is where the problems start

OpenStudy (anonymous):

\[ \frac{\partial (fu)}{\partial x} = \frac{ \partial f}{\partial x} u + f\frac{\partial u}{\partial x}\] etc....

OpenStudy (anonymous):

those are equal what u just wrote above?

OpenStudy (anonymous):

That's only the first term...

OpenStudy (anonymous):

oh i think i just clicked what ur saying

OpenStudy (anonymous):

okay..so the entire left side is now expanded

OpenStudy (anonymous):

okay....i think i see where this is going now

OpenStudy (anonymous):

if i expand the right side it's gonna come out the same way huh?

OpenStudy (anonymous):

Yeah but it would probably be better if you just grouped the terms on the left. \[f \frac{\partial u}{\partial x} + f\frac{\partial v}{\partial y} + f\frac{\partial w}{\partial z} = f(\vec{\nabla} \cdot \vec{V})\] and so forth..

OpenStudy (anonymous):

very good

OpenStudy (anonymous):

thank you soo soo much